OpenStudy (grace111):

Please help = medal and fan . (1) - I am thinking of a 2-digit number that is a multiple of 8 . The digits add up to six . What is the number ? (2) - It takes one and a half minutes to swim a length . How many length can I swim in 12 minutes ?

OpenStudy (anonymous):

(1) 24. 8 x 3 = 24 2 + 4 = 6 (2) 8 lengths. 12/1.5 = 8

OpenStudy (grace111):

thank you soo much !